If the dollar falls by 20% against the euro and rises by 10% against the yen, which of the following values for European and Japanese trade with the United States are consistent with a 10% increase in the effective exchange of the United States?

A) Europe: 33%; Japan: 66% B) Europe: 66%; Japan: 33% C) Europe: 50%; Japan: 50% D) None of these values is consistent with this increase.

The answer is D, none of the above. But why? I thought it is B because 20(0.66) -10(0.33) = 9.9 (according to the textbook 's definition, a fall in $ value means a rise in exchange rate)
Answer #1

Let the value of European trade be E and Japanese be J

we have, E + J = 1 ....................(1)

and, -20E + 10J = 10................(2)

Thus, -20(1-J) + 10J = 10

=> -10J = -10

=> J =1 => E =0

Thus, the entire tarde is with Japan

Thus, (d) is the correct answer as all other options have E>0